OpenStudy (anonymous):

Boyle's Law says that the volume of a gas varies inversely with the pressure. When the volume of a certain gas is 5L , the pressure is 452 kPa (kilopascals). What is the volume when the pressure is 565 kPa?

jimthompson5910 (jim_thompson5910):

If the volume varies inversely with pressure then V*P = k where V = volume, P = pressure So 5*452 = k and V*565 = k which means 5*452 = V*565 2260 = V*565 2260/565 = V 4 = V V = 4 Which means that the volume of the gas is 4 liters